Fill in the missing items for the following inventories:

A B C
Beginning balance $52,200 $28,900 $68,500
Ending balance 46,600 25,700 54,200
Transferred in 52,400
Transferred out 93,100 170,600
Answer #1

Answer

  • Equation used:
    Beginning balance + Transferred in – Transferred out = Ending balance

A

B

C

Beginning balance

$52,200

$28,900

$68,500

Ending balance

$46,600

$25,700

$54,200

Transferred in

$52,400

$89,900

$156,300

Transferred out

$58,000

$93,100

$170,600

  • Working

A

B

C

Beginning balance

52200

28900

68500

Ending balance

46600

25700

54200

Transferred in

52400

=25700+93100-28900

=54200+170600-68500

Transferred out

=52200+52400-46600

93100

170600